OpenStudy (anonymous):

What is the solution set of |x – 6| = 10?

jimthompson5910 (jim_thompson5910):

If |x – 6| = 10 then x-6 = 10 or x-6 = -10

jimthompson5910 (jim_thompson5910):

Solve each equation to find your answers

OpenStudy (anonymous):

would it be 16 and -4

jimthompson5910 (jim_thompson5910):

you got it